# Understanding Epistemic Language with a Bayesian Theory of Mind

### v1

@@ 3.2.2 Joint Inference of Goals and Beliefs
3.2.3 Evaluating Epistemic Sentences
By inferring the agent’s belief history $b_{0:T}$, we can now compute the probability $\mathbf{Pr}(A,\phi)$ assigned to a formula $\phi$ at time $t$ as the weighted sum of truth values under the agent’s belief $b_{t}$:
However, observers do not have direct access to the true $s_{t}$ or $b_{t}$, only probabilistic inferences about them. As such, we model human evaluation of a sentence $\varphi$ as a quantitative probabilistic judgment given their posterior inferences:
4 Experiments
To evaluate our model on a diverse set of belief statements, we conducted a two-part experiment. We first recruited participants to write English sentences describing the current and initial beliefs of a player character as it navigated a gridworld puzzle that required finding keys hidden in boxes, collecting a dataset of epistemic language (Table 2). Next, we asked two groups of participants to rate how likely these sentences were to be true given the player’s observed behavior, with one group rating sentences about the player’s current beliefs, and the other group rating sentences about initial beliefs. These ratings were collected at multiple points over the course of the player’s trajectory. We then compared these ratings against the inferences produced by our model.
4.1 Scenario Construction
@@ 4.2 Collecting Epistemic Language
In the first part of the experiment, we recruited 42 US participants via Prolific (mean age: 36.02, SD: 10.1; 16 women, 26 men). Following a tutorial, participants watched 10 scenario animations, with each stopping before the player reached their goal and all relevant keys were revealed. Participants were then asked to write at least two sentences about the player’s likely beliefs at the end of the scenario (current beliefs), and another two sentences about the player’s beliefs at the start of the scenario (initial beliefs). To ensure that these sentences focused on beliefs about the environment, we instructed participants to describe the player’s beliefs about the contents of the boxes. We excluded 8 participants for failing to follow these instructions, and about one third of remaining sentences. [^1] See Appendix for exclusion and post-processing steps. Interestingly, some excluded sentences described the player’s beliefs about what keys were needed (e.g. “The player thinks that they need to open box 3 to get to the hexagonal gem.”). We save the study of these sentences for future work. This process left us with 241 (228) statements about current (initial) beliefs, which were then annotated with factors by two experimenters. Table 2 shows examples illustrating the diversity of language we collected.
@@ 4.3.1 Statement Selection
To ensure that the belief statements evaluated by our participants were (i) diverse and (ii) rated enough times to ensure sufficient statistical power (88% power at Cohen’s $d=0.8$), we selected a set of 5 statements per scenario (3 plausible, 2 implausible) from our full dataset of collected statements. The plausible statements were chosen by repeatedly sampling sets of 3 statements out of all those written for a particular scenario, then selecting the set that scored highest on a diversity metric derived from the factors in Table 2 (see Appendix for the definition of this metric). We then manually added 2 more statements that were originally written for other scenarios, and which we evaluated to be implausible descriptions of the target scenario. Participants were shown 2 of these 5 statements at random in each scenario.
4.4 Model Fitting and Evaluation
We evaluated our LaBToM model on all 20 scenarios, producing normalized likelihood scores for the 5 current and 5 initial belief statements per scenario. We then computed the correlation coefficient $r$ between these scores and average human ratings. We fit our model parameters to maximize $r$, fitting the belief thresholds $\Theta:=(\theta_{\textsf{believes}},\theta_{\textsf{could}}...)$ via coordinate ascent, and the inverse temperature $\beta$ of our action model via grid search. This produced the fitted values for $\Theta$ shown in Table 1, and $\beta=2^{3/2}$. For the set of possible initial agent beliefs $\mathcal{B}_{0}$, we fixed the number of belief particles to $k=3$ to ensure the tractability of exact Bayesian inference.
@@ 4.5 Baselines
To assess the import of a sufficiently rich theory of mind for epistemic language understanding, we evaluated several ablations of our LabToM model which made simplified assumptions about the agent’s beliefs or planning abilities. We also evaluated two state-of-the-art multi-modal LLMs, thereby testing the degree to which grounded evaluation of epistemic language can be achieved with sufficient scale and finetuning:
@@
Full (ours) 0.81 (0.01) 0.81 (0.01) 0.80 (0.01)
Non-Planning 0.47 (0.01) 0.60 (0.02) 0.20 (0.01)
True Belief 0.12 (0.01) 0.09 (0.01) 0.13 (0.02)
GPT-4o (OpenAI et al., 2023)
@@
Gemini 1.5 Pro (Gemini Team et al., 2024)
@@
Table 3: Model correlations with human ratings of epistemic language. LaBToM correlates strongly with humans, whereas multimodal LLMs struggle.
Non-Planning.
Multi-modal LLMs
We use GPT-4o (text & image input, gpt-4o-2024-05-13) and Gemini 1.5 Pro (text & image or video input) as multi-modal LLM baselines, providing them the same instructions as human participants. Each baseline was run 3 times with a temperature of 1.0. In addition to providing an image or video of the scenario showing the actions up to each judgment point, we used the following prompting methods:
  • Plans: The prompt plainly describes the agent’s actions over time (e.g. the player moves right five times), and also describes the agent’s observations (e.g. the player opens box 1 and finds a red key inside).
  • Narratives: The prompt contains a rich narrative of the agent’s actions, providing key information about the scene (e.g. which gems are locked behind doors, which keys are visible), while also describing the agent’s movements in relation to relevant objects (e.g. the player moves right five times, going past box 1 towards box 2).
  • Few-Shot Prompting: After describing the plan or narrative, we provide the LLM with human ratings for the 4 other belief statements tied that scenario, before querying its rating for the target statement.
5 Results
Figure 2: Human correlation plots for LaBToM and GPT-4o. LaBToM provides a much stronger qualitative fit with human ratings compared to the best GPT-4o baseline, which fails to use the full scale.
@@ 5 Results
As Table 3 shows, we find that that our full model produces statement scores that correlate highly with human ratings ($r$ = $0.81$) across both current and initial belief conditions (a per-factor breakdown can be found in the Appendix). Plotting average human judgments against LaBToM inferences (Figure 2, Row 1), we also find a strong qualitative fit: LaBToM generally assigns high or low ratings to sentences when humans do, using the ends of the scale as appropriate. In contrast, our ablated models do poorly, either because they fail to track how the player updates their beliefs (True Belief, Current $r$ = $0.09$), or fail to infer the player’s initial beliefs from their actions towards instrumental subgoals (Non-Planning, Initial $r$ = $0.13$).
@@
The multi-modal LLM baselines also perform less well than LaBToM despite extensive prompting, with the strongest LLM baseline (GPT-4o with images, narratives, and few-shot prompting) achieving human correlations of only 0.59 and 0.41 respectively. Increasing information in the prompt improves performance, although the Gemini models perform poorly regardless, and do not improve even when provided full video of the scenarios. We find that LLM performance is lower when evaluating initial belief sentences, suggesting that they are better at tracking how agents’ current beliefs change in response to observations, but worse at inferring past beliefs.
@@
We analyze just one scenario (Figure 3a) in detail due to space constraints. In this scenario, the player first walks upwards away from box 3 (Judgment Point 1), leading both humans and LaBToM to assign scores of greater than 0.5 to “The player believes that box 3 is empty” and “The player initially believed that box 3 was empty”. In contrast, the modal sentence “The player believes that box 3 may contain a red key” is rated as less likely. This is because if the player did believe strongly that box 3 had a (red) key, then acting rationally, it should have looked inside that box, but this does not occur.
After the player opens box 2 and finds it empty, then walks back down towards box 3 (Judgment Point 3), both humans and our model decrease their confidence in the statement “The player believes that box 3 is empty”, as expected. They also decrease confidence in “The player initially believed that box 3 was empty”, but less sharply. This is because there are at least two possibilities consistent with the player’s actions: They could have firmly believed that box 3 was empty, or they could have just believed that box 3 was less likely to contain the relevant key than box 2, without all-out believing that it was empty. This ability to understand beliefs about relative likelihood is made explicit by how ratings change over time for “The player believes that box 3 is more likely to contain a red key than box 1.” Consistent with the principle of rational action, both humans and LaBToM assign a high score to this sentence once the player goes back to box 3, forgoing box 1.
Across a range of other settings (Figure 3b–c), our model largely captures fine-grained changes in how people evaluate a variety of epistemic expressions, including modal sentences, ignorance claims, and expressions of uncertainty. Unlike prior BToM models that lack language-like belief representations (Baker et al., 2017), LaBToM also distinguishes observer and agent uncertainty, assigning high ratings to claims that the agent is uncertain (Figure 3c, S5 Current), and vice versa. In contrast, the best performing LLM baseline (GPT-4o, I+Na+FS) often fails to adjust its ratings in the same direction as humans, while giving high ratings to implausible statements. We discuss the step-by-step results in greater detail in the Appendix, alongside cases where our model comes apart from human judgments.
@@
To investigate how our model generalizes to a larger set of epistemic expressions, we performed the in-context vs. out-of-context likelihood comparison described in Section 4.4 for our full dataset of 469 sentences. We tested the full LaBToM model, ablations and the best applicable LLM baseline from Table 3 (GPT-4o, I+Na). Results are shown in Table 4. For both current and initial belief statements, we find that LaBToM assigns significantly higher scores when a statement is evaluated in-context vs. out-of-context, correctly classifying the context about 70% of the time by assigning a strictly higher in-context score. As we examine in the Appendix, many statements that are not correctly classified turn out to be plausible in either context, resulting in equal or close-to-equal scores. The ablated models and GPT-4o perform significantly worse than LaBToM, especially for initial beliefs, where all baselines perform worse-than-chance.
@@ 6 Discussion
Our experiments show that, similar to humans, our LaBToM model is able to coherently interpret and adjust its evaluations of natural language statements about agents’ beliefs, whereas state-of-the-art multimodal LLMs struggle with this task. We also find that LaBToM largely distinguishes in vs. out-of-context sentence usage on a large and diverse set of crowd-sourced epistemic language, indicating the generalizability of our approach.
That said, our model is not without limitations. As we discuss at greater length in the Appendix, LaBToM’s outputs depart from human judgments in several interesting ways, suggesting the need to account for (i) contextual adaptation of probability thresholds Schuster and Degen (2020), (ii) the role of justification in human’s intuitive evaluation of knowledge claims Alston (1989), and (iii) bounded human reasoning about logical implications Smets and Solaki (2018). LaBToM is also an ideal observer model that does not scale readily to large belief spaces, leaving open how humans tractably infer and evaluate claims about others’ beliefs Van Rooij (2008), perhaps by focusing on occurent beliefs (Bartlett, 2018) that are relevant to others’ goals. Finally, LaBToM is a model of how people interpret epistemic language, but full understanding also includes the ability to produce such language. This could potentially be achieved by inverting the ELoT module of our model, using it to translate salient inferences about an agent’s beliefs into natural language. By extending our model in this way, we stand to gain an even richer account of what it means to understand epistemic language.
Acknowledgements
We thank our colleagues Cedegao Zhang, Brian Leahy, and Megan Wei for helpful discussions during the development of this project. We would also like to thank Jacob Andreas and Cedegao Zhang for valuable feedback on the paper.
This work was funded in part by the DARPA Machine Common Sense, AFOSR, and ONR Science of AI programs, along with the MIT-IBM Watson AI Lab, the Aphorism Foundation and the Siegel Family Foundation. Tan Zhi-Xuan is supported by an Open Phil AI Fellowship.

### v2

@@ 4.2 Collecting Epistemic Language
In the first part of the experiment, we recruited 42 US participants via Prolific (mean age: 36.02, SD: 10.1; 16 women, 26 men). Following a tutorial, participants watched 10 scenario animations, with each stopping before the player reached their goal and all relevant keys were revealed. Participants were then asked to write at least two sentences about the player’s likely beliefs at the end of the scenario (current beliefs), and another two sentences about the player’s beliefs at the start of the scenario (initial beliefs). To ensure that these sentences focused on beliefs about the environment, we instructed participants to describe the player’s beliefs about the contents of the boxes. We excluded 8 participants for failing to follow these instructions, and about one third of remaining sentences (see Appendix). This process left us with 241 (228) statements about current (initial) beliefs, which were then annotated with factors by two experimenters. Table 2 illustrates the diversity of language we collected.
@@ 4.3.1 Statement Selection
To ensure that the belief statements evaluated by our participants were (i) diverse and (ii) rated enough times to ensure sufficient statistical power (88% power at Cohen’s $d=0.8$), we selected 5 statements per scenario (3 plausible, 2 implausible) from our full dataset of statements. The plausible statements were chosen by sampling many sets of 3 statements out of all those written for a scenario, then selecting the set that scored highest on a diversity metric derived from the factors in Table 2 (see Appendix for details). We then manually added 2 more statements that were originally written for other scenarios, and which we evaluated to be implausible descriptions of the target scenario. Participants were shown 2 of these 5 statements at random in each scenario.
4.4 ELoT Parser Evaluation
We evaluated the suitability of ELoT formulas as representations of epistemic sentences by investigating (i) the impact of using ELoT as the translation target (instead of the lowered form, which lacks epistemic operators besides $\mathbf{Pr}$) and (ii) the value of the associated ELoT grammar $G_{\mathcal{E}}$ in constraining LLM outputs.
To do so, we created gold translations of the subset of statements selected for human evaluation (125 statements) into ELoT formulas and their lowered forms. We then compared the translations produced by grammar-constrained SMC decoding ($n_{\sigma}$$=$$10$ samples) of LLaMa 3.1 8B against these gold translations in terms of both strict equality and approximate semantic equivalence (i.e. cases of reasonable alternative translations of the original English statement). As baselines, we used unconstrained self-consistency sampling Wang et al. (2023) with LLaMa 3.1 8B and the instruction-tuned Gemini Flash 8B, sampling $n_{\sigma}$$=$$10$ times and taking the majority answer. All methods used the same prompt format (see Appendix), with 36 example translations.
4.5 LaBToM Fitting and Evaluation
We evaluated our LaBToM model on all 20 scenarios, producing normalized likelihood scores for the 5 current and 5 initial belief statements per scenario. We then computed Pearson’s $r$ between these scores and average human ratings. We fit our model parameters to maximize $r$, fitting the belief thresholds $\Theta:=(\theta_{\textsf{believes}},\theta_{\textsf{could}}...)$ via coordinate ascent, and the inverse temperature $\beta$ via grid search (see Appendix for robustness analyses). This produced the fitted values for $\Theta$ in Table 1, and $\beta=2^{3/2}$. For the set of possible initial agent beliefs $\mathcal{B}_{0}$, we fixed the number of belief particles to $k=3$ to ensure tractable exact inference.
@@ 4.6 Baselines
To assess the import of a sufficiently rich theory of mind for epistemic language understanding, we evaluated several ablations of our model under simplified assumptions about the agent’s beliefs or planning abilities. We also evaluated two state-of-the-art multi-modal LLMs, thereby testing the degree to which grounded evaluation of epistemic language can be achieved with sufficient scale:
@@
Non-Planning.
Multi-modal LLMs.
We use GPT-4o (text & image input, gpt-4o-2024-05-13) and Gemini 1.5 Pro (text & image or video input) as multi-modal LLM baselines, providing them the same instructions as human participants (see Appendix for a unimodal baseline). Each baseline was run 3 times with a temperature of 1.0. In addition to providing an image or video of the scenario showing the actions up to each judgment point, we used the following prompting methods:
Plans: The prompt plainly describes the agent’s actions over time (e.g. the player moves right five times), and also describes the agent’s observations (e.g. the player opens box 1 and finds a red key).
Narratives: A rich narrative of agent behavior is included in the prompt, providing key information about the scene (e.g. which gems are locked behind doors, which keys are visible), while also describing the agent’s movements in relation to relevant objects (e.g. the player moves right five times, going past box 1 towards box 2).
Few-Shot Prompting: After describing the plan or narrative, we provide the LLM with human ratings for the 4 other statements tied that scenario, before querying its rating for the target statement.
5 Results
Parser Translation Accuracy (s.e.)
ELoT Lowered
Exact Equiv. Exact Equiv.
Grammar-Constrained SMC Loula et al. (2025)
LLaMa 3.1 8B
0.81
0.91
0.34
0.42
Unconstrained Sampling + Majority Vote
LLaMa 3.1 8B
0.65
0.69
0.32
0.36
Gemini Flash 8B
0.69
0.80
0.60
0.68
Table 3: Translation accuracy for ELoT vs. lowered formulas as the translation target, compared across sampling methods and LLMs in terms of exact equality (Exact) or semantic equivalence (Equiv.). ELoT serves as a better target and constraint for translation.
Parser Example
English The player believes that box 3 is empty.
ELoT Gold believes(player, formula(empty(box3)))
Grammar-Cons. believes(player, formula(empty(box3)))
Unconstrained believes(player, empty(box3)) (Syntax Error)
Lowered Gold >=(prob_of(player, empty(box3)), threshold(believes))
Grammar-Cons. and(>=(prob_of(player, empty(box3)), threshold(believes)), empty(box3)) (Semantic Error)
Unconstrained >=(prob_of(player, empty(box3)), believes) (Syntax Error)
Table 4: Example translations from English to ELoT and lowered formulas using LLaMa 3.1 8B.
ELoT is a superior translation target and guide for (grammar-constrained) LLM parsing.
We report translation accuracy in Table 3. In confirmation of the idea that our ELoT formalism captures the semantics of belief sentences at the right level of granularity, we find that translating English into ELoT is up to 2.4 times as accurate as translating into the lowered form, even though they have the same expressive capacity. ELoT also increases the benefit of grammar-constrained SMC decoding, leading to an improvement of 0.20 in equivalence accuracy for LLaMa 3.1 8B. In contrast, using the lowered grammar leads to only an 0.06 improvement over unconstrained generation. We show example translations and errors for each representation in Table 4. These findings highlight the importance of “language-of-thought” representations that are aligned with the structure of natural language, in line with work tying language use to the structure and acquisition of compositional concepts Wong et al. (2022).
@@ 5 Results
Full (ours) 0.76 (0.01) 0.78 (0.01) 0.72 (0.01)
Non-Planning 0.40 (0.01) 0.58 (0.01) 0.07 (0.01)
True Belief 0.10 (0.01) 0.09 (0.01) 0.09 (0.02)
GPT-4o OpenAI et al. (2023)
@@
Gemini 1.5 Pro Gemini Team et al. (2024)
@@
Table 5: Model correlations with human ratings of epistemic language. LaBToM correlates strongly with humans, whereas multimodal LLMs struggle to do so.
Overall, grammar-constrained SMC decoding into ELoT with LLaMa 3.1 8B performs the best, achieving 91% equivalence with our gold translations. Upon inspection (see Appendix), we find that unconstrained sampling often leads to syntax errors not corrected for by majority voting. Gemini Flash 8B also exhibits low sample diversity typical of instruction tuning, such that majority voting gives no improvement. In comparison, grammar-constrained SMC produces not just accurate translations, but a distribution over reasonable translations, since SMC performs posterior sampling rather than optimization Loula et al. (2025); Lew et al. (2023b). In future work, this could be used to more precisely model the interpretation of ambiguous epistemic language.
Figure 2: Human correlation plots for LaBToM and GPT-4o. Black dots are from statements not accurately translated to ELoT. LaBToM provides a much stronger qualitative fit with human ratings compared to the best GPT-4o baseline, which fails to use the full scale.
@@
As Table 5 shows, our full model produces statement scores that correlate highly with human ratings ($r$ = $0.76$) across both current and initial belief conditions. In the Appendix, we show results using the gold ELoT translations ($r$ = $0.81$), and a per-factor breakdown. Plotting average human judgments against LaBToM inferences (Figure 2, Row 1), we also find a strong qualitative fit: Factoring out incorrect ELoT translations (black dots), LaBToM generally assigns high or low ratings to sentences when humans do, using the ends of the scale as appropriate. In contrast, our ablated models do poorly, either because they fail to track how the player updates their beliefs (True Belief, Current $r$ = $0.09$), or fail to infer the player’s initial beliefs (Non-Planning, Initial $r$ = $0.07$). These findings hold even when the parameters of our full model are adversarially optimized (see Appendix).
@@
The multi-modal LLM baselines also perform less well than LaBToM despite extensive prompting, with the strongest LLM baseline (GPT-4o with images, narratives, and few-shot prompting) achieving human correlations of only 0.59 and 0.41 respectively. Increasing information in the prompt improves performance, although the Gemini models do poorly regardless, even with full scenario videos. We find that LLM performance is lower on initial belief sentences, suggesting that they are better at tracking how agents’ current beliefs change, but worse at inferring past beliefs.
@@
We analyze just one scenario (Figure 3a) in detail. Here, the player first walks upwards away from box 3 (Judgment Point 1), leading both humans and LaBToM to assign scores of greater than 0.5 to “The player believes that box 3 is empty” and “The player initially believed that box 3 was empty”. In contrast, the modal sentence “The player believes that box 3 may contain a red key” is rated lower. This is because if the player did believe that box 3 might have the (red) key, then it is more likely that they would have looked in box 3. However, this does not occur.
After the player opens box 2 and finds it empty, then walks back down towards box 3 (Judgment Point 3), both humans and our model decrease their confidence in the statement “The player believes that box 3 is empty”. They also decrease confidence in “The player initially believed that box 3 was empty”, but less sharply. This is because there are at least two possibilities consistent with the player’s actions: They could have firmly believed that box 3 was empty, or they could have just believed that box 3 was less likely to contain the relevant key than box 2, without all-out believing that it was empty. This ability to understand beliefs about relative likelihood is made explicit by how ratings change over time for “The player believes that box 3 is more likely to contain a red key than box 1.” Consistent with the principle of rational action, both humans and LaBToM assign a high score to this sentence once the player goes back to box 3, forgoing box 1.
Across a range of other settings (Figure 3b–c), our model largely captures fine-grained changes in how people evaluate a variety of epistemic expressions, including modal sentences, ignorance claims, and expressions of uncertainty. Unlike prior BToM models that lack language-like belief representations Baker et al. (2017), LaBToM also distinguishes observer and agent uncertainty, assigning high ratings to claims that the agent is uncertain (Figure 3c, S5 Cur.), and vice versa. In contrast, the best LLM baseline (GPT-4o, I+Na+FS) often fails to adjust its ratings in the same direction as humans, while rating implausible statements highly. We discuss these results more in the Appendix, alongside cases where our model comes apart from humans.
@@
To investigate how our model generalizes to a larger set of epistemic expressions, we performed the in-context vs. out-of-context likelihood comparison described in Section 4.5 for our full dataset of 469 sentences. We tested the full LaBToM model, ablations and the best applicable LLM baseline from Table 5 (GPT-4o, I+Na). Results are shown in Table 6. For both current and initial belief statements, we find that LaBToM assigns significantly higher scores when a statement is evaluated in-context vs. out-of-context, correctly classifying the context about 70% of the time by assigning a strictly higher in-context score. On closer inspection, many statements that are incorrectly classified turn out to be plausible in either context, resulting in equal or close-to-equal scores. The ablated models and GPT-4o perform significantly worse than LaBToM, especially for initial beliefs.
@@ 6 Discussion
Our experiments show that, similar to humans, our LaBToM model is able to coherently interpret and adjust its evaluations of natural language statements about agents’ beliefs, whereas state-of-the-art multimodal LLMs struggle with this task. This ability is mediated by our ELoT representation and semantic parser: Without a compositional representation of epistemic concepts that is aligned with the structure of natural epistemic language, we find that translation accuracy drops significantly, even when using a grammar-constrained LLM parser. We also find that LaBToM largely distinguishes in vs. out-of-context sentence usage on a large and diverse set of crowd-sourced epistemic language, indicating the generalizability of our approach.
That said, our model is not without limitations. As we discuss at greater length in the Appendix, LaBToM’s outputs depart from human judgments in several interesting ways, suggesting the need to account for (i) contextual adaptation of probability thresholds via pragmatic reasoning Rudin (2016); Schuster and Degen (2020), (ii) the role of justification in human’s intuitive evaluation of knowledge claims Alston (1989), and (iii) bounded human reasoning about logical implications Smets and Solaki (2018).
LaBToM is also an ideal observer model that does not scale readily to large belief spaces, leaving open how humans tractably infer and evaluate claims about others’ beliefs Van Rooij (2008), perhaps by focusing on occurent beliefs Bartlett (2018) that are relevant to others’ goals. Finally, LaBToM is a model of how people interpret epistemic language, but full understanding also includes the ability to produce such language. This could potentially be achieved by inverting the ELoT module of our model, using it to translate salient or conversationally-relevant inferences about an agent’s beliefs into natural language. By extending our model in this way, we stand to gain an even richer account of what it means to understand epistemic language.
Acknowledgments
We thank our colleagues Cedegao Zhang, Brian Leahy, and Megan Wei for helpful discussions in the development of this project, and Ben LeBrun for his help in applying grammar-constrained SMC decoding to our setting. We also thank Jacob Andreas, Cedegao Zhang, Liam Bright, Daniel Lassiter, our reviewers, and our action editor for their valuable feedback on this paper.
This work was funded in part by the DARPA Machine Common Sense, AFOSR, and ONR Science of AI programs, along with the MIT-IBM Watson AI Lab, the Siegel Family Foundation, and an anonymous donor. Tan Zhi-Xuan is supported by an Open Philanthropy AI Fellowship.